1) Suppose a perfectly competitive labour market has a demand curve of LD = 120 − 2w and a supply curve of LS = 8w, where w is the wage rate is dollars and L is the quantity of labour in person-hours
a) What are the equilibrium values of the wage and employment?
b) Suppose the government imposed a minimum wage of $14 per hour. Now what are the equilibrium values of the wage and employment?

c) Repeat part (a), assuming now that the market is a monopsony.
d) Repeat part (b), assuming now that the market is a monopsony.

e) Does the imposition of the minimum wage decrease employment here under perfect competition? What about under monopsony? Give a brief intuitive explanation for your answer and why it may be different under the two different market structures.
1
Expert's answer
2020-04-13T09:44:52-0400

a) LD=LS

120-2w=8w

w=12, LS=96

b) LD=120-28=92, w=14

c)The total expenditure on labor, or w(L) · L (where w(L) is the inverse supply of labor), is L2/8.

Thus, the marginal expenditure on labor is L/4. Setting marginal expenditure equal to marginal

benefit (the inverse of the original labor demand) gives L/4 = 60 − L/2, or L = 80. The wage

rate is found where L = 80 on the labor supply curve, which is at w = 10

Employment will decrease

d) the values ​​will not change, since monopsony will not increase the rate of care payment, and workers will not go anywhere, so the indicators will remain at the same level

e) The introduction of a minimum wage in the conditions of perfect competition will reduce employment, as firms will need to legally pay for this minimum wage level, plus the balance of the spent funds. It is economically disadvantageous for firms in conditions of perfect competition.

In monopsony, with the introduction of a minimum wage, the level of employment may remain the same or increase. Since the monopsony firm is one on the labor market and workers have no where to go, as they agree with the minimum wage.